what is the acceleration of sphere rolling down a plane inclined at an angle 30° to horizontal?
1
Expert's answer
2020-12-28T12:56:04-0500

"ma=mg\sin\alpha-F_f"


"I\epsilon=r\cdot F_f\to (2/5)mr^2a/r=rF_f\to F_f=(2/5)ma"


"ma=mg\sin\alpha-(2/5)ma \to"


"a=(5/7)g\sin\alpha=(5/7)\cdot9.8\cdot\sin30°=3.5(m/s^2)"